OpenStudy (anonymous):

Given the differential equation (x^2+2x-1)y''-2(x+1)y'+2y=0 Show that the equation has a linear polynomial and a quadratic polynomial as solutions.

OpenStudy (dumbcow):

For linear: start with y = mx+b then y' = m y'' = 0 plug these into diff equ and simplify, if all the "x" terms cancel out then "y" can be linear under certain conditions do same thing for quadratic: start with y = ax^2 +bx +c y' = 2ax +b y'' = 2a

OpenStudy (anonymous):

thank you very much. easy enough when explained. :-). so for linear I got m=b, and for quadratic i got c=a+b. I am then asked to come up with two linearly independent solutions of equation and give general solution. so all i need to do is just let y=mx+b, and y=ax^2+bx+c, and set m=c=1, or any constant, and I get y=x+1 and y=x^2+x+2, and use both of those for my general solution.... does it matter what I set the constant = too.... could I of used y=2x+2 and y=2x^2+2x+4 instead?

OpenStudy (dumbcow):

no doesn't matter what constant you pick....for general solution , just recognize fact that if a=0 in quadratic solution you end up with linear solution y = ax^2 +bx + (a+b) a=0 y = bx + b thus general solution is: y = ax^2 + bx+ (a+b) for unique solution, you would need 2 initial values to determine constants (a,b)